OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle?

OpenStudy (anonymous):

THIS is where i draw it out on graph paper

OpenStudy (espex):

You know that the perimeter (p) is equal to twice the length (l) plus twice the width (w). So p = 2l + 2w \(\rightarrow\) 52 = 2l + 2w You also know that the length is 8 inches longer than the width, l = w + 8 or l - w = 8 Giving you a system of equations, pick your method and solve.